OpenStudy (anonymous):

Fifty-four of the 90 books on a shelf are history books. What percent of the books are history books?

OpenStudy (anonymous):

Divide 54 by 90 and then multiply the resulting decimal by 100 to convert to a percent. 54 / 90 = 0.60, then 0.60 * 100 = 60%

OpenStudy (anonymous):

Thankies. :)

OpenStudy (anonymous):

You could also set up a proportion: 54/90 = x/100 and solve for x via cross products.